A 1129kg automobile is speeding at 25m/s. What is the resultant force required to stop
1
Expert's answer
2020-06-24T11:21:34-0400

According to Newton's second law, the resultant force required to stop the car is mass times acceleration. The acceleration depends on the time required to stop:


"F=ma=mm\frac{v}{t}=1129\cdot\frac{25}{t}=\frac{28225}{t}\text{ N}."